OpenStudy (anonymous):

A small child is at a candy store, eager to spend his $1 allowance on candy. His two favorite candies are priced at $0.05 and $0.10, respectively. If the child is intent on spending his entire allowance on candy, which of the following represents a standard form equation that describes this situation?

OpenStudy (anonymous):

A.0.05x+0.10y=1.00 B.0.10x-0.05y=1.00 c.5x+10y=100.00 d.x+2y=20.00

OpenStudy (anonymous):

help me plzzz

OpenStudy (anonymous):

I think its 2th .

OpenStudy (anonymous):

wats 2th?

OpenStudy (anonymous):

The : B.0.10x-0.05y=1.00

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

i wanna see the other persons answer first \

OpenStudy (calculusxy):

What this question is asking for you to do is finding what expression best explains the situation. You can ask yourself the question of "How many candies of $0.05 and $0.10 can the person buy?" So you also should know that if a variable is right next to the number, then it is to be multiplied. The best answer that you can find should be A. You multiply $0.05 and $0.10 and find the different variations of finding how many candies can the person buy.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

can you plzz help me with more :(

OpenStudy (anonymous):

Try all : 1 .: It s fall because the .050.10 are not the numbers of the candies .! 3.:The child didnt buy 5 or 10 candies .! 4.:It isnt the problem !!!

OpenStudy (calculusxy):

@andieboo1995 I will be more than happy to help you for your further questions.

OpenStudy (anonymous):

@calculusxy: You re realy right !:)

OpenStudy (anonymous):

thank you sooooo much :)

OpenStudy (anonymous):

What do you want ?!

OpenStudy (anonymous):

sorry i didnt finish it

OpenStudy (calculusxy):

Okay. My fault.

OpenStudy (anonymous):

Use substitution to determine which of the following points is a solution to the standard form equation below. 5x-2y=10 -1,5 1,5 0,-5 0,5

OpenStudy (anonymous):

@andieboo1995 :Do you want x & y ?

OpenStudy (anonymous):

they didnt say

OpenStudy (anonymous):

they just said which points

OpenStudy (anonymous):

It is A!!

OpenStudy (anonymous):

OK ! The points are x & y friend !:)

OpenStudy (anonymous):

-1,5 @Jonathan1997 ?

OpenStudy (anonymous):

Now look :

OpenStudy (nirmalnema):

i think the c option is correct

OpenStudy (anonymous):

Get all of them in this .

OpenStudy (anonymous):

Sorry, it is C

OpenStudy (anonymous):

i have another one A high school basketball team has a budget specifically for towels and extra basketballs. A towel costs $6 and a basketball costs $30. Assume the budget is $300 and that 25 towels are needed for the team for the entire season. Use a standard form equation to determine how many basketballs may be purchased by the team. A. 5 B. 15 C. 30 D. 275

OpenStudy (nirmalnema):

y=5x-10 / 2 if u put x= 0 you will get -10/2 = -5 so (0,-5)

OpenStudy (anonymous):

oh okay so is is C EVERYONE!!!!

OpenStudy (anonymous):

A-5.

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

can you explain y @Jonathan1997

OpenStudy (anonymous):

i have more

OpenStudy (anonymous):

30x=150

OpenStudy (anonymous):

It s for basketball costs !

OpenStudy (anonymous):

so its C @E.ali

OpenStudy (calculusxy):

$25(6)+$30x=$300 And now solve for x.

OpenStudy (anonymous):

Yes. $6 * 25 towels = 150. 150 divided by 300 = 150. 150 divided by 300 = 5

OpenStudy (anonymous):

@ andieboo1995 :yeah !

OpenStudy (calculusxy):

So as according to Jonathan1997, the high school basketball team needs at least 5 basketballs to fit into the centre of the budget.

OpenStudy (anonymous):

so its 5 @calculusxy

OpenStudy (anonymous):

Rewrite the equation below in standard form.

OpenStudy (anonymous):

What do you want ?!:)

OpenStudy (anonymous):

it in standard form :?

OpenStudy (anonymous):

standard form for what ?!:)

OpenStudy (anonymous):

the equation above

OpenStudy (calculusxy):

Did you find the answer for the equation give above?